# references.bib

### v1

@@ 4.2.1 The XX spin chain
$r(u)=\left(\begin{array}[]{cccc}1&0&0&0\\ 0&\tan(u)&\sec(u)&0\\ 0&\sec(u)&\tan(u)&0\\ 0&0&0&1\\ \end{array}\right).$ (4.22)
After constructing the quantum gate $R(u)$, differently from the previous two cases, by solving the right boundary Yang-Baxter equation (2.9), we obtain two independent types of solutions
@@
For the solution of the type f, as for the interacting case, $k_{i}$ and $\tilde{k}_{j}$ are solutions of the one layer spin chain and take the form (4.16).
The solution of type nf are the most interesting and will allow us to answer positively to the question posed at the beginning. In fact, for this case, considering the boundary Yang-Baxter equation in a larger space where the $R$-matrix is the tensor product of two $r$-matrices, allows for a non-factorized form of the solutions. We solved the boundary YBE (2.9) as explained in sec. 3.2 by starting from the most general ansatz for the $K^{R}$-matrix with all the 16 elements and we obtained that all the solutions of the right boundary Yang-Baxter equation are of $8$ vertex type, explicitly
$K^{R}(u)=\left(\begin{array}[]{cccc}\frac{\kappa_{1,1}-\kappa_{3,3}+2\cot u}{q(u)}-1&0&0&\frac{\kappa_{1,4}}{q(u)}\\ 0&1&\frac{\kappa_{2,3}}{q(u)}&0\\ 0&\frac{\kappa_{3,2}}{q(u)}&\frac{\kappa_{3,3}}{q(u)}+1&0\\ \frac{\kappa_{4,1}}{q(u)}&0&0&\frac{2\cot u-\kappa_{1,1}}{q(u)}-1\\ \end{array}\right)$ (4.23)
@@
$q(u)=\frac{1}{4\cot u}\left(\kappa_{2,3}\kappa_{3,2}+\kappa_{1,1}\left(\kappa_{3,3}-\kappa_{1,1}\right)-\kappa_{1,4}\kappa_{4,1}-2\kappa_{3,3}\cot u+4\cot^{2}u\right).$ (4.24)
@@
It is clear that this $K^{R}(u)$ matrix cannot be written as the tensor product of two $2\times 2$ $k$-matrices.
To obtain the solution of the $K^{L}(u)$, we use one of the three (equivalent) automorphisms of [Sklyanin:1988yz], summarized in section 2.1.2. This model has crossing unitarity symmetry [^15] Contrary to the interacting model, here the crossing unitarity property is preserved for the $R$-matrix in the enlarged space. The reason is that here $r(u-2\eta)=r(u+2\eta)$. and $\eta=\pi/2$.
@@
$K^{L,(i)}(u)=\Big{(}K^{R,(i)}\Big{)}^{t}(-u-\pi/2),\,\,\,\,\,i\in\{\text{f},\text{nf}\}.$ (4.25)
For this model, we are allowed to build 4 different type of transfer matrices. In fact, there are two different families of $K^{R}$ matrices that solve the right boundary Yang-Baxter eq. (2.9), which we labelled $K^{R,\text{(f)}}$ and $K^{R,\text{(nf)}}$ and two that solve the eq. (2.10) ($K^{L,\text{(f)}}$, $K^{L,\text{(nf)}}$).
By following Sklyanin’s proof [Sklyanin:1988yz] of the commutativity between $[t(u),t(v)]=0$ for (2.15), one can similarly prove that,
$t^{(i,j)}(u,\mathbf{w})\,=\,\text{tr}_{0}\Big{(}K_{0}^{L,(i)}(u)T(u,\mathbf{w})K_{0}^{R,(j)}(u)\hat{T}(u,\mathbf{w})\Big{)},$ (4.26)
$[t^{(i,j)}(u,\mathbf{w}),t^{(i,j)}(v,\mathbf{w})]=0,\,\,\,\,\,\,\,\,\,\,\forall\,\,\,i,j\in\{\text{f},\,\text{nf}\}\,.$ (4.27)
The defition (4.26) for the transfer matrix $t(u)$ includes 4 transfer matrices. This type of construction was also considered in [shiroishi1997integrable] for the Hubbard model.
@@
For this type of circuit, the space-reflection symmetry is broken, while integrability is preserved. This type of circuit is equivalent to a finite free fermionic model with an interacting impurity. By fixing the parameters $\kappa_{1,4}=\kappa_{4,1}=0$ in (4.23) and considering the unitarity conditions given in sec. 4.3, we obtain unitary dynamics with an interacting impurity with $U(1)$ symmetry. It would perhaps be interesting to analyze the transport properties of this model.
@@ 4.2.2 The XX spin chain in magnetic field
We consider the case of the XX spin chain in a magnetic field with strength $h$. This model is also referred to as the XXh model. The $r$-matrix is [gomez1996quantum],
$r(u)=\left(\begin{array}[]{cccc}1-e^{i(u+2\psi)}&0&0&0\\ 0&e^{i\psi}\left(1-e^{iu}\right)&-2ie^{\frac{1}{2}i(u+2\psi)}\sin\psi&0\\ 0&-2ie^{\frac{1}{2}i(u+2\psi)}\sin\psi&e^{i\psi}\left(1-e^{iu}\right)&0\\ 0&0&0&e^{iu}-e^{2i\psi}\\ \end{array}\right),$ (4.32)
@@
Due to the presence of the magnetic field, the solution of type f has less free parameters as compared to the previous case. Only a diagonal solution is allowed,
$k^{R}(u)=\left(\begin{array}[]{cc}\frac{2}{1-\kappa_{1,1}\tan\left(\frac{u}{2}\right)}-1&0\\ 0&1\\ \end{array}\right).$ (4.35)
@@
$K^{R}(u)=\left(\begin{array}[]{cccc}\frac{\cot\left(\frac{u}{2}\right)-\kappa_{4,4}}{g(u)}-1&0&0&\frac{\kappa_{1,4}}{g(u)}\\ 0&1&0&0\\ 0&0&\frac{\kappa_{3,3}}{g(u)}+1&0\\ \frac{\kappa_{4,1}}{g(u)}&0&0&\frac{\kappa_{4,4}-\kappa_{3,3}+\cot\left(\frac{u}{2}\right)}{g(u)}-1\\ \end{array}\right),$ (4.36)
@@
$K^{L,(i)}(u)=\Big{(}K^{R,(i)}\Big{)}^{t}(-u-\pi),\,\,\,\,\,i\in\{\text{f},\text{nf}\}.$ (4.37)
Since also for the case with the magnetic field, we obtained two families of $K$-matrices, we can construct four different transfer matrices commuting as (4.26) and (4.27).
@@
$r(u)=\left(\begin{array}[]{cccc}1&0&0&0\\ 0&\frac{2u}{2u+1}&\frac{1}{2u+1}&0\\ 0&\frac{1}{2u+1}&\frac{2u}{2u+1}&0\\ 0&0&0&\frac{1-2u}{2u+1}\\ \end{array}\right).$ (4.42)
@@
$K^{R,\text{(f)}}_{ij}=k^{R}_{i}(u)\tilde{k}^{R}_{j}(-u),\,\,\,\,\,\text{where}\,\,\,\,\,k^{R}(u)=\left(\begin{array}[]{cc}\frac{2}{1-u\,\kappa_{1,1}}-1&0\\ 0&1\\ \end{array}\right),$ (4.45)
@@
$K^{R,\text{(nf)}}(u)=\left(\begin{array}[]{cccc}\frac{8-\kappa_{4,4}}{m(u)}-1&0&0&\frac{\kappa_{1,4}}{m(u)}\\ 0&1&0&0\\ 0&0&\frac{\kappa_{3,3}}{m(u)}+1&0\\ \frac{\kappa_{4,1}}{m(u)}&0&0&\frac{\kappa_{3,3}-\kappa_{4,4}-2/u}{m(u)}-1\\ \end{array}\right),$ (4.50)
@@
For this model, the crossing unitarity property is broken even at the level of $r(u)$. We obtain the solutions $K^{L,\text{(f)}}$ and $K^{L,\text{(nf)}}$ from the automorphism (2.20),
$K^{L,\text{(f)}}_{i,j}(u)=k^{L}_{i}(u)\tilde{k}^{L}_{j}(-u),\,\,\,\,\,\text{where}\,\,\,k^{L}(u)=\left(\begin{array}[]{cc}\frac{\kappa_{1}(2u-1)+1}{\kappa_{1}(2u+1)-1}&0\\ 0&1\\ \end{array}\right),$ (4.53)
$K^{L,\text{(nf)}}(u)=\left(\begin{array}[]{cccc}\frac{-8\left(\kappa_{3,3}-2u\kappa_{4,4}+2\right)+h(-2u)-h(-1)}{h(1-2u)}&0&0&\frac{16u\kappa_{1,4}}{h(1-2u)}\\ 0&1&0&0\\ 0&0&\frac{h(2u+1)}{h(1-2u)}&0\\ \frac{16u\kappa_{4,1}}{h(1-2u)}&0&0&\frac{-8\left(\kappa_{3,3}+2u\kappa_{4,4}+2\right)+h(2u)-h(-1)}{h(1-2u)}\\ \end{array}\right),$ (4.58)
where $h(u)=u^{2}\kappa_{4,4}^{2}+u^{2}\kappa_{1,4}\kappa_{4,1}-u\kappa_{3,3}\left(u\kappa_{4,4}+4\right)-16$.
Also for this case, we can construct $4$ different commuting transfer matrices as (4.26) and (4.27).
@@ 4.3 Unitarity of the $U$ and $K$ gates
In this paragraph, we discuss the unitarity property of the quantum gates that we used to build the quantum circuit. In particular, in section 2.2.2, we obtain that to build the quantum circuit for the case with open boundary condition, the main ingredients are the quantum gate $U$ given in (2.23) and $K^{R}$, solution of the boundary reflection algebra (2.9) and $\bar{K}^{L}$, connected to the solution of the reflection algebra by (2.32).
@@ 5 Continuous time dynamics
In this section, we focus on the continuous time dynamics. In sec. 5.1, we obtain the continuous dynamics as a limit of the discrete one. We motivate our choice of the factorized form of the quantum gate. In sec. 5.2, we discuss the connection between the boundary terms in the conserved charges and their Lindbladian structure.
@@ 5.1 Continuous time dynamics as the limit of discrete one
$\check{R}_{i,j}(u)=\mathbb{I}+u\,\partial_{u}\check{R}_{i,j}(u)|_{u=0}+\dots=\mathbb{I}+u\,q^{(2)}_{i,j}+\dots,$ (5.1)
@@
$U_{i,j}=\check{R}_{i,j}(2\kappa)$ (5.2)
and the discrete dynamical evolution for two steps is generated by the $M$ operator given in (2.26) for the case with periodic boundary condition and (2.33) for the open one.
To obtain the continuous time evolution, [vanicat2018integrable2], we set $u=-i\,\Delta t$ and $\Delta t=t/n$ and we apply $n$ times the two step discrete evolution $M$, with $n$ being large. In this way, we obtain the continuous time evolution
$\lim_{n\to\infty}M^{n}=\exp(-i\,t\,\mathbb{Q}_{2}),$ (5.3)
@@
We clarify the reason of our choice of the quantum gate $R(u)=r(u)\otimes r(-u)$ by adding the indices explicitly (3.1) and taking the limit $u\to 0$. The subscripts in the operator identify the space on the original spin chain where the operators are acting non-trivially. The charge associated to this $R$-matrix is
$\partial_{u}r_{ik}(u)r_{jl}(-u)|_{u\to 0}=q^{(2)}_{ij,kl}=h_{ik}-h_{jl}=h_{ik}-h^{t}_{jl},$ (5.5)
@@
At this stage, a step back is necessary. Initially, our motivation for exploring this specific problem was to determine whether the property of a model’s steady state being integrable in a boundary-driven setup could also extend to the Yang-Baxter integrability of the spin chain. Specifically, we considered the research from over a decade ago on constructing non-equilibrium steady state (NESS) density operators for boundary-driven, locally interacting quantum chains, where the driving is applied through Markovian dissipation channels (Lindblad evolution) at the chain’s boundaries, [prosen2011open, prosen2011exact, prosen2015matrix, ilievski2014exact, ilievski2014quantum, prosen-exterior-lindblad, vznidarivc2010exact].
@@
$\text{End}(\mathcal{H})\to\mathcal{H}\otimes\mathcal{H}^{*}.$ (5.7)
@@
$\mathcal{L}=\sum_{J}\mathcal{L}_{J,J^{\prime}}\equiv i\sum_{j}(h_{2j,2j+2}-h_{2j+1,2j+3}^{t}),$ (5.8)
where with the sum over $J$ we are considering the neighbouring super-sites in the vectorized space $J=(2j,2j+1),J^{\prime}=(2j+2,2j+3)$. This coincides, up to a normalization factor, with the expression given in (5.5).
In the previous sections, by constructing the reflection algebra, we answered to the question whether we are allowed to add some non-factorizable terms to the boundary of the spin chain, such that the Yang-Baxter integrability of the model is preserved. We obtained that, if we consider an interacting model in the bulk of the spin chain, meaning the XXX or the XXZ spin chain, there are no interesting (non-factorizable) boundary term that preserves integrability. However, for the case of the XX spin chain, both with and without magnetic field, we obtained non-trivial solutions of the right (2.9) and left reflection algebra (2.10).
In [prosen2011open, prosen2011exact], it is shown that for the XXZ chain with Lindbladian spin source/sink on the left/right boundary, the NESS is proven to be exactly solvable in terms of a matrix product ansatz. This, however, does not coincide with the full Yang-Baxter integrability of the spin chain. In fact, the only case where we found a non-factorizable solution is the case where in the bulk we have the XX model. We now focus on this type of solution to understand if it is possible to write the boundary terms in the Lindbladian form.

### v2

@@ 4.2.1 The XX spin chain
$r(u)=\left(\begin{array}[]{cccc}1&0&0&0\\ 0&\tan(u)&\sec(u)&0\\ 0&\sec(u)&\tan(u)&0\\ 0&0&0&1\\ \end{array}\right).$ (61)
After constructing the quantum gate $R(u)$, differently from the previous two cases, by solving the right boundary Yang-Baxter equation (10), we obtain two independent types of solutions
@@
For the solution of the type f, as for the interacting case, $k_{i}$ and $\tilde{k}_{j}$ are solutions of the one layer spin chain and take the form (55).
The solution of type nf are the most interesting and will allow us to answer positively to the question posed at the beginning. In fact, for this case, considering the boundary Yang-Baxter equation in a larger space where the $R$-matrix is the tensor product of two $r$-matrices, allows for a non-factorized form of the solutions. We solved the boundary YBE (10) as explained in sec. 3.2 by starting from the most general ansatz for the $K^{R}$-matrix with all the 16 elements and we obtained that all the solutions of the right boundary Yang-Baxter equation are of $8$ vertex type, explicitly
$K^{R}(u)=\left(\begin{array}[]{cccc}\frac{\kappa_{1,1}-\kappa_{3,3}+2\cot u}{q(u)}-1&0&0&\frac{\kappa_{1,4}}{q(u)}\\ 0&1&\frac{\kappa_{2,3}}{q(u)}&0\\ 0&\frac{\kappa_{3,2}}{q(u)}&\frac{\kappa_{3,3}}{q(u)}+1&0\\ \frac{\kappa_{4,1}}{q(u)}&0&0&\frac{2\cot u-\kappa_{1,1}}{q(u)}-1\\ \end{array}\right)$ (62)
@@
$q(u)=\frac{1}{4\cot u}\left(\kappa_{2,3}\kappa_{3,2}+\kappa_{1,1}\left(\kappa_{3,3}-\kappa_{1,1}\right)-\kappa_{1,4}\kappa_{4,1}-2\kappa_{3,3}\cot u+4\cot^{2}u\right).$ (63)
@@
It is clear that this $K^{R}(u)$ matrix cannot be written as the tensor product of two $2\times 2$ $k$-matrices, hence it is not factorized.
To obtain the solution of the $K^{L}(u)$, we use one of the three (equivalent) automorphisms of [3], summarized in section 2.1.2. This model has crossing unitarity symmetry [^16] Contrary to the interacting model, here the crossing unitarity property is preserved for the $R$-matrix in the enlarged space. The reason is that here $r(u-2\eta)=r(u+2\eta)$. and $\eta=\pi/2$.
@@
$K^{L,(i)}(u)=\Big{(}K^{R,(i)}\Big{)}^{t}(-u-\pi/2),\,\,\,\,\,i\in\{\text{f},\text{nf}\}.$ (64)
For this model, we are allowed to build 4 different type of transfer matrices. In fact, there are two different families of $K^{R}$ matrices that solve the right boundary Yang-Baxter eq. (10), which we labelled $K^{R,\text{(f)}}$ and $K^{R,\text{(nf)}}$ and two that solve the eq. (11) ($K^{L,\text{(f)}}$, $K^{L,\text{(nf)}}$).
By following Sklyanin’s proof [3] of the commutativity between $[t(u),t(v)]=0$ for (16), one can similarly prove that,
$t^{(i,j)}(u,\mathbf{w})\,=\,\text{tr}_{0}\Big{(}K_{0}^{L,(i)}(u)T(u,\mathbf{w})K_{0}^{R,(j)}(u)\hat{T}(u,\mathbf{w})\Big{)},$ (65)
$[t^{(i,j)}(u,\mathbf{w}),t^{(i,j)}(v,\mathbf{w})]=0,\,\,\,\,\,\,\,\,\,\,\forall\,\,\,i,j\in\{\text{f},\,\text{nf}\}\,.$ (66)
The defition (65) for the transfer matrix $t(u)$ includes 4 transfer matrices. This type of construction was also considered in [43] for the Hubbard model.
@@
For this type of circuit, the space-reflection symmetry is broken, while integrability is preserved. This type of circuit is equivalent to a finite free fermionic model with an interacting impurity. By fixing the parameters $\kappa_{1,4}=\kappa_{4,1}=0$ in (62) and considering the unitarity conditions given in sec. 4.3, we obtain unitary dynamics with an interacting impurity with $U(1)$ symmetry. It would perhaps be interesting to analyze the transport properties of this model.
@@ 4.2.2 The XX spin chain in magnetic field
We consider the case of the XX spin chain in a magnetic field with strength $h$. This model is also referred to as the XXh model. The $r$-matrix is [44],
$r(u)=\left(\begin{array}[]{cccc}1-e^{i(u+2\psi)}&0&0&0\\ 0&e^{i\psi}\left(1-e^{iu}\right)&-2ie^{\frac{1}{2}i(u+2\psi)}\sin\psi&0\\ 0&-2ie^{\frac{1}{2}i(u+2\psi)}\sin\psi&e^{i\psi}\left(1-e^{iu}\right)&0\\ 0&0&0&e^{iu}-e^{2i\psi}\\ \end{array}\right),$ (71)
@@
Due to the presence of the magnetic field, the solutions of type f and nf have less free parameters as compared to the previous cases.
For the type f, only a diagonal solution is allowed,
$k^{R}(u)=\left(\begin{array}[]{cc}\frac{2}{1-\kappa_{1,1}\tan\left(\frac{u}{2}\right)}-1&0\\ 0&1\\ \end{array}\right).$ (74)
@@
$K^{R}(u)=\left(\begin{array}[]{cccc}\frac{\cot\left(\frac{u}{2}\right)-\kappa_{4,4}}{g(u)}-1&0&0&\frac{\kappa_{1,4}}{g(u)}\\ 0&1&0&0\\ 0&0&\frac{\kappa_{3,3}}{g(u)}+1&0\\ \frac{\kappa_{4,1}}{g(u)}&0&0&\frac{\kappa_{4,4}-\kappa_{3,3}+\cot\left(\frac{u}{2}\right)}{g(u)}-1\\ \end{array}\right),$ (75)
@@
$K^{L,(i)}(u)=\Big{(}K^{R,(i)}\Big{)}^{t}(-u-\pi),\,\,\,\,\,i\in\{\text{f},\text{nf}\}.$ (76)
Since also for the case with the magnetic field, we obtained two families of $K$-matrices, we can construct four different commuting transfer matrices as (65) and (66).
@@
$r(u)=\left(\begin{array}[]{cccc}1&0&0&0\\ 0&\frac{2u}{2u+1}&\frac{1}{2u+1}&0\\ 0&\frac{1}{2u+1}&\frac{2u}{2u+1}&0\\ 0&0&0&\frac{1-2u}{2u+1}\\ \end{array}\right).$ (81)
@@
$K^{R,\text{(f)}}_{ij}=k^{R}_{i}(u)\tilde{k}^{R}_{j}(-u),\,\,\,\,\,\text{where}\,\,\,\,\,k^{R}(u)=\left(\begin{array}[]{cc}\frac{2}{1-u\,\kappa_{1,1}}-1&0\\ 0&1\\ \end{array}\right),$ (84)
@@
$K^{R,\text{(nf)}}(u)=\left(\begin{array}[]{cccc}\frac{2/u-\kappa_{4,4}}{m(u)}-1&0&0&\frac{\kappa_{1,4}}{m(u)}\\ 0&1&0&0\\ 0&0&\frac{\kappa_{3,3}}{m(u)}+1&0\\ \frac{\kappa_{4,1}}{m(u)}&0&0&\frac{\kappa_{4,4}-\kappa_{3,3}-2/u}{m(u)}-1\\ \end{array}\right),$ (89)
@@
For this model, the crossing unitarity property is broken even at the level of $r(u)$. We obtain the solutions $K^{L,\text{(f)}}$ and $K^{L,\text{(nf)}}$ from the automorphism (21),
$K^{L,\text{(f)}}_{i,j}(u)=k^{L}_{i}(u)\tilde{k}^{L}_{j}(-u),\,\,\,\,\,\text{where}\,\,\,k^{L}(u)=\left(\begin{array}[]{cc}\frac{\kappa_{1}(2u-1)+1}{\kappa_{1}(2u+1)-1}&0\\ 0&1\\ \end{array}\right),$ (92)
$K^{L,\text{(nf)}}(u)=z(u)\left(\begin{array}[]{cccc}\frac{-8\left(\kappa_{3,3}-2u\kappa_{4,4}+2\right)+h(-2u)-h(-1)}{h(1-2u)}&0&0&\frac{16u\kappa_{1,4}}{h(1-2u)}\\ 0&1&0&0\\ 0&0&\frac{h(2u+1)}{h(1-2u)}&0\\ \frac{16u\kappa_{4,1}}{h(1-2u)}&0&0&\frac{-8\left(\kappa_{3,3}+2u\kappa_{4,4}+2\right)+h(2u)-h(-1)}{h(1-2u)}\\ \end{array}\right),$ (97)
where $h(u)=u^{2}\kappa_{4,4}^{2}+u^{2}\kappa_{1,4}\kappa_{4,1}-u\kappa_{3,3}\left(u\kappa_{4,4}+4\right)-16$ and $z(u)=\frac{\left(1-16u^{2}\right)\,h(1-2u)}{64u^{2}\left(6\left(u\kappa_{3,3}+2\right)+h(u)\right)}$.
Also for this case, we can construct $4$ different commuting transfer matrices as (65) and (66).
@@ 4.3 Unitarity of the $U$ and $K$ gates
In this paragraph, we discuss the unitarity property of the quantum gates that we used to build the quantum circuit. In particular, in section 2.2.2, we obtain that to build the quantum circuit for the case with open boundary condition, the main ingredients are the quantum gate $U$ given in (24) and $K^{R}$, solution of the boundary reflection algebra (10) and $\bar{K}^{L}$, connected to the solution of the reflection algebra by (33).
@@ 5 Continuous time dynamics
In this section, we focus on the continuous time dynamics. In sec. 5.1, we obtain the continuous dynamics as a limit of the discrete one. We motivate our choice of the factorized form $R(u)=r(u)\otimes r(-u)$ of the quantum gate. In sec. 5.2, we discuss the connection between the boundary terms in the conserved charges and their Lindbladian structure.
@@ 5.1 Continuous time dynamics as the limit of discrete one
$\check{R}_{i,j}(u)=\mathbb{I}+u\,\partial_{u}\check{R}_{i,j}(u)|_{u=0}+\dots=\mathbb{I}+u\,q^{(2)}_{i,j}+\dots,$ (98)
@@
$U_{i,j}=\check{R}_{i,j}(2\kappa)$ (99)
and the discrete dynamical evolution for two steps is generated by the $M$ operator given in (27)-(28) for the case with periodic boundary condition and (34)-(35) for the open one.
To obtain the continuous time evolution, [9], we set ${2\kappa}=-i\,\Delta t$ and $\Delta t=t/n$ and we apply $n$ times the two step discrete evolution $M$, with $n$ being large. In this way, we obtain the continuous time evolution
$\lim_{n\to\infty}M^{n}=\exp(-i\,t\,\mathbb{Q}_{2}),$ (100)
@@
We clarify the reason of our choice of the quantum gate $R(u)=r(u)\otimes r(-u)$ by adding the indices explicitly (37) and taking the limit $u\to 0$. The subscripts in the operator identify the space on the original spin 1/2 chain where the operators are acting non-trivially. The charge associated to this $R$-matrix is
$\partial_{u}r_{ik}(u)r_{jl}(-u)|_{u\to 0}=q^{(2)}_{ij,kl}=h_{ik}-h_{jl}=h_{ik}-h^{t}_{jl},$ (102)
@@
At this stage, a step back is necessary. Initially, our motivation for exploring this specific problem was to determine whether the property of a model’s steady state being integrable in a boundary-driven setup could also extend to the Yang-Baxter integrability of the spin chain with open boundary condition. Specifically, we considered the research from over a decade ago on constructing non-equilibrium steady state (NESS) density operators for boundary-driven, locally interacting quantum chains, where the driving is applied through Markovian dissipation channels (Lindblad evolution) at the chain’s boundaries, [23, 24, 25, 26, 27, 28, 29].
@@
$\text{End}(\mathcal{H})\to\mathcal{H}\otimes\mathcal{H}^{*}.$ (104)
@@
$\mathcal{L}=\sum_{J}\mathcal{L}_{J,J^{\prime}}\equiv i\sum_{j=0}(h_{2j+1,2j+3}-h_{2j+2,2j+4}^{t}),$ (105)
where with the sum over $J$ we are considering the neighbouring super-sites in the vectorized space $J=(2j+1,2j+2),\,J^{\prime}=(2j+3,2j+4)$. This coincides, up to a normalization factor, with the expression given in (102).
In the previous sections, by constructing the reflection algebra, we answered to the question whether we are allowed to add some non-factorizable terms to the boundary of the spin chain, such that the Yang-Baxter integrability of the model is preserved. We obtained that, if we consider an interacting model in the bulk of the spin chain, meaning the XXX or the XXZ spin chain, there are no interesting (non-factorizable) boundary term that preserves integrability. However, for the case of the XX spin chain, both with and without magnetic field, we obtained non-trivial solutions of the right (10) and left reflection algebra (11).
In [23, 24], it is shown that for the XXZ chain with Lindbladian spin source/sink on the left/right boundary, the NESS is proven to be exactly solvable in terms of a matrix product ansatz. This, however, does not coincide with the full Yang-Baxter integrability of the spin chain. In fact, the only case where we found a non-factorizable solution is the case where in the bulk we have the XX model. We now focus on this type of solution to understand if it is possible to write the boundary terms in the Lindbladian form.